OpenStudy (anonymous):

Solve the system. x2 + y2 = 4 y - x = 2

OpenStudy (angela210793):

x2 + y2 = 4 y - x = 2-->y=2+x so we can replace 2+x instead of y in the 1st eq.---->x^2+(2+x)^2=4 x^2+4+4x+x^2=4 2X^2+4x+4=4 x^2+2x+2=1 x^2+2x+1=0

OpenStudy (anonymous):

rewrite the second one as y = x+2 and substitute into the first one to get \[x^2+(x+2)^2=4\]

OpenStudy (anonymous):

what angela said!

OpenStudy (angela210793):

u may continue Satellite ....my fingers r tired :):):)

OpenStudy (anonymous):

square the second equation on both sides and then add bith equatios we get x=0 and y=2